# Gromov’s Compactness Theorem for the Intrinsic Timed Hausdorff Distance

### v1

@@ 4. Applications of Addresses
In this section we restate Gromov’s Compactness Theorem and an Arzela-Ascoli Theorem with addresses and state an open question.
@@ 4.1. Gromov’s Compactness Theorem with Addresses
If $(X_{j},d_{j})$ are a sequence of compact metric spaces that are equibounded as in (4) and equicompact as in (5) then a subsequence converges in the Gromov-Hausdorff sense to a compact metric space $(\bar{X}_{\infty},d_{\infty},\tau_{\infty})$. In fact, there exist distance preserving Fréchet maps,
(149) $\kappa_{j}:X_{j}\to Z\subset\ell^{\infty},$
@@
In addition there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(151) ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}\textrm{ and }{\mathcal{I}}^{\infty}:{\mathcal{A}}\to\bar{X}_{\infty}$
such that we have uniform convergence as $j\to\infty$,
(152) $\sup_{\alpha\in{\mathcal{A}}}\|\kappa_{j}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0.$
In particular,
@@
This theorem follows exactly as in the proof of our Theorem 3.1 simply by removing all references to the time function. For completeness of exposition, we prove it as a trivial corollary of our theorem:
@@
4.2. Arzela-Ascoli Theorem with Addresses
Theorem 4.2 (Arzela-Ascoli-with-Addresses).
@@ 4.2. Arzela-Ascoli Theorem with Addresses
(155) $(X_{j},d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(X_{\infty},d_{\infty})$
with a compact limit space, and given uniformly bounded $K$-Lipschitz functions for some $K>0$,
(156) $F_{j}:X_{j}\to[0,F_{max}],$
@@
(157) $F_{\infty}:X_{\infty}\to Y_{\infty}$
in the following sense. There exist distance preserving Fréchet maps,
(158) $\kappa_{j}:X_{j}\to Z\subset\ell^{\infty},$
@@
In addition there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(160) ${\mathcal{I}}^{j}:{\mathcal{A}}\to X_{j}\textrm{ and }{\mathcal{I}}^{\infty}:{\mathcal{A}}\to\bar{X}_{\infty}$
@@
(161) $\sup_{\alpha\in{\mathcal{A}}}\|\kappa_{j}({\mathcal{I}}^{j}(\alpha))-\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\|_{\infty}\to 0,$
@@
4.3. Arzela-Ascoli Conjecture with Addresses
It would be interesting to apply the address method to more precisely describe the following stronger Arzela-Ascoli Theorem:
@@ 4.3. Arzela-Ascoli Conjecture with Addresses
Given a pair of Gromov-Hausdorff converging sequences of compact metric space:
(164) $X_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}X_{\infty}\textrm{ and }Y_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}Y_{\infty}$
and given uniformly bounded $K$-Lipschitz functions for some $K>0$,
(165) $F_{j}:X_{j}\to Y_{j},$
then a subsequence converges to a bounded $K$-Lipschitz function
(166) $F_{\infty}:X_{\infty}\to Y_{\infty}.$

### v2

@@ 4. Applications of Addresses
In this final section we restate Gromov’s Compactness Theorem and an Arzelà–Ascoli Theorem in terms of addresses and pose an open question about a more general version of Arzelà-Ascoli.
@@ 4.1. Gromov’s Compactness Theorem with Addresses
If $(X_{j},d_{j})$ is a sequence of compact metric spaces that are equibounded as in (9) and equicompact as in (10) then a subsequence converges in the Gromov–Hausdorff sense to a compact metric space $(\bar{X}_{\infty},d_{\infty})$. In fact, there exist distance preserving Fréchet maps,
(154) $\kappa_{j}\colon X_{j}\to Z\subset\ell^{\infty},$
@@
In addition, there is an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(156) ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}\quad\textrm{ and }\quad{\mathcal{I}}^{\infty}\colon{\mathcal{A}}\to\bar{X}_{\infty}$
such that
(157) $\sup_{\alpha\in{\mathcal{A}}}d_{\ell^{\infty}}\Big(\kappa_{j}({\mathcal{I}}^{j}(\alpha)),\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\Big)\to 0$
as $j\to\infty$. In particular,
@@
This theorem follows exactly as in the proof of our Theorem 3.1 simply by removing all references to the time function. For completeness of exposition, we prove it as a trivial corollary of our theorem.
@@
4.2. Arzelà–Ascoli Theorem with Addresses
Theorem 4.2 (Arzelà–Ascoli-with-Addresses).
@@ 4.2. Arzelà–Ascoli Theorem with Addresses
(160) $(X_{j},d_{j})\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}(X_{\infty},d_{\infty}),$
and uniformly bounded $K$-Lipschitz functions for some $K>0$,
(161) $F_{j}\colon X_{j}\to[0,F_{max}],$
@@
(162) $F_{\infty}\colon X_{\infty}\to[0,F_{max}]$
in the following sense: There exist distance preserving Fréchet maps,
(163) $\kappa_{j}\colon X_{j}\to Z\subset\ell^{\infty},$
@@
In addition there exist an uncountable index set of addresses, $\mathcal{A}$, and surjective index maps,
(165) ${\mathcal{I}}^{j}\colon{\mathcal{A}}\to X_{j}\quad\textrm{ and }\quad{\mathcal{I}}^{\infty}\colon{\mathcal{A}}\to\bar{X}_{\infty}$
@@
(166) $\sup_{\alpha\in{\mathcal{A}}}d_{\ell^{\infty}}\Big(\kappa_{j}({\mathcal{I}}^{j}(\alpha)),\kappa_{\infty}({\mathcal{I}}^{\infty}(\alpha))\Big)\to 0,$
@@
4.3. Arzelà–Ascoli Conjecture with Addresses
It would be interesting to apply the address method to more precisely describe the following stronger Arzelà–Ascoli Theorem:
@@ 4.3. Arzelà–Ascoli Conjecture with Addresses
Given a pair of Gromov–Hausdorff converging sequences of compact metric spaces:
(169) $X_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}X_{\infty}\quad\textrm{ and }\quad Y_{j}\stackrel{{\scriptstyle\textrm{GH}}}{{\longrightarrow}}Y_{\infty}$
and given uniformly bounded Lipschitz $K$ functions for some $K>0$,
(170) $F_{j}\colon X_{j}\to Y_{j},$
then a subsequence converges to a bounded Lipschitz $K$ function
(171) $F_{\infty}\colon X_{\infty}\to Y_{\infty}.$